Q. Given the rate of the third order reaction:
A+B+C→P is 0.05 Ms-1
If the [A] = 0.05 M, [B] = 0.01M, and [C] = 0.25M. What is the third order rate constant?

Q. What are the units of the rate constant for a second-order reaction?

Q. Given the first order reaction is completed 90% in 30 minutes at 298 K. Calculate the rate
1
Expert's answer
2020-07-16T07:00:58-0400

Given reaction

(1) A+B+C --------- P

Rate of reaction =0.05MS-1

R = K.(A)1.(B)1.(C)1

0.05 =K. (0.05) . (0.01) . (0.25)

K = 0.05/(0.05) .(0.01). (0.25)

K =10000/25 = 400 M-2. S-1

Rate of constant = 400 M2.S-1

(2) Unit of rate constant for a second order reaction M-1.S-1 OR L.mol-1.s-1

(3) For first order reaction

Rate constant K = 2.303/30 log100/10

K = 20303/30 LOG10 = 2.303/30 min-1 = 0.076min-1